Find the product using suitable properties
1005 x 168

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given 1005 × 168
= (1000 + 5) × 168
= 1000 × 168 + 5 × 168 (using distributive property)
= 168000 + 840
= 168840

Thanks